You are dealt one card from a 52 card deck. Find the probability that you are not dealt a card from 6-9?
Answer by fractalier(6550)

There are sixteen cards from 6 to 9. That means there are 52-16 = 36 cards that are not 6 thru 9...so
The probability is 36/52 = 9/13